OpenStudy (anonymous):

How many grams of water must be added to 2 grams of a 65% iodine solution in order to produce a 42% iodine solution?

OpenStudy (anonymous):

x+2=y 0x+(2)(.65)=(.42)(y) 1.3=(.42)(y) y=3.095 y=total solution 3.095-2=1.095 1.095 grams of water (I think)